A motorcycle is moving with a velocity of 10ms-1, seeing a roadblock at the front, he slowed down until it stops with 7 seconds. What is the deceleration of the motorcycle?


1
Expert's answer
2021-02-05T04:29:44-0500
"a=\dfrac{v-v_0}{t},""a=\dfrac{0-10\ \dfrac{m}{s}}{7\ s}=-1.43\ \dfrac{m}{s^2}."